OpenStudy (anonymous):

EMERGENCY: NEED HELP Thank You Two blocks with masses m1 = 3.9 kg and m2 = 6.7 kg are connected by a string that hangs over a pulley of mass M = 2.2 kg and radius R = 0.11 m as shown above. The string does not slip. Assuming the system starts from rest, use energy principle find the speed of m2 after it has fallen by 0.4 m. Treat the pulley as a disk.

OpenStudy (anonymous):

emergency?

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

must of pated the whole thing

OpenStudy (anonymous):

pasted

OpenStudy (anonymous):

\[1/2 I^2 \omega=m gh\]

OpenStudy (anonymous):

you know a diagram would really help

OpenStudy (anonymous):



OpenStudy (anonymous):

here's the diagram

OpenStudy (anonymous):

obviously, it will turn toward mass with bigger mass

OpenStudy (anonymous):

net mass 6.7-3.9=2.8 2.8*9.8=27.44 newton

OpenStudy (anonymous):

\[1/2 I^2 \omega=m gh\] \[1/2 I^2 \omega=27.44(.4)\] you can find I and plug it in. then solve for \[\omega\] once you find \[\omega\], divide by r

OpenStudy (anonymous):

is mass for both or only the heavier one

OpenStudy (anonymous):

this 6.7-3.9=2.8

OpenStudy (anonymous):

ok I'll solve it add see if its right

OpenStudy (anonymous):

yeah he incorporates integrals into the equations

OpenStudy (anonymous):

Is I= mr^2/2

OpenStudy (anonymous):

if so, is m for both masses?

OpenStudy (anonymous):

Is I= mr^2/2 yes if so, is m for both masses? net mass

OpenStudy (anonymous):

ok thank you, im gonna solve for the acce..

OpenStudy (anonymous):

\[1/2 I^2 \omega=27.44(.4)\] once you find I , plug it here and solve for omega

OpenStudy (anonymous):

I'm not quite getting realistic numbers

OpenStudy (anonymous):

I= 1/2 m r^2 = 1/2 (2.2)(.11)^2 I=0.01331

OpenStudy (anonymous):

i get 6097.77 for w

OpenStudy (anonymous):

wrong thing square \[1/2 I \omega^2=27.44(.4)\] \[1/2 (0.01331) \omega^2=27.44(.4)\] w~~40.6114

OpenStudy (anonymous):

oh tats why

OpenStudy (anonymous):

divide that by by radius 40.6114/.11m

OpenStudy (anonymous):

.11 meter that is

OpenStudy (anonymous):

369.19?

OpenStudy (anonymous):

that doesn't sound right

OpenStudy (anonymous):

and if I multiply I get 4.467

OpenStudy (anonymous):

which is also the wrong answer for some reason

OpenStudy (anonymous):

is 63.63 right answer?

OpenStudy (anonymous):

I'll check

OpenStudy (anonymous):

Don't think its right

OpenStudy (anonymous):

this question is frustrating

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

try 421.084

OpenStudy (anonymous):

did you try it?